# Video Anomaly Detection in 10 Years: A Survey and Outlook

### v1

@@ V-B2 Deep Feature Extractors
While each feature representation has its strengths and weaknesses as empirically investigated by the SOTA VAD approaches [6, 26, 35, 50]. The end-to-end deep feature learning paradigms are the recent trends and these approaches have advanced the VAD performance. We provide the following future directions in terms of the feature representations to improve detection, accuracy, and efficiency.
@@ V-C1 Supervised Approaches
In supervised learning, algorithms are trained on datasets where frames are pre-classified as either ’normal’ or ’anomalous.’ However, supervised methods are less common due to the scarcity of datasets with detailed annotations at the frame level for the training set.
@@ V-C2 Self-supervised Approaches
In self-supervised approaches, algorithms generate their supervision from the input data, bypassing the need for external labels provided by humans. This paradigm is particularly valuable for anomaly detection, where labeled anomalous data is scarce.
In the work by [9], researchers propose a method based on self-supervised and multi-task learning at the object level. Self-supervision involves training a 3D CNN on several proxy tasks that do not require labeled anomaly data. These tasks include determining the direction of object movement (arrow of time), identifying motion irregularities by comparing objects in consecutive versus intermittent frames, and reconstructing object appearances based on their preceding and succeeding frames. By learning from the video data itself what normal object behavior looks like, the model becomes adept at detecting anomalies when deviations from this learned behavior occur. This methodology enables effective anomaly detection even in the absence of explicit labels.
@@ V-C3 Weakly Supervised Approaches
In weakly supervised approaches, obtaining precise frame-level annotations for anomalies in long video sequences can be challenging and time-consuming. Instead, annotators may label “snippets” or short segments within the video where anomalies are observed, serving as weak labels for training the model.
In their work, [11] introduced a pioneering multiple instance learning (MIL) model, the first to adopt the notion of weakly labeled training videos. Here, normal and anomalous videos are considered negative and positive bags, respectively, with video segments serving as instances in MIL. These bags are then processed through feature extractors to glean spatiotemporal features before being directed to a fully connected network to produce the final output. The output anomaly score ranges from 0 to 1, with the goal of increasing the model’s output anomaly score for abnormal segments while decreasing it for normal segments.
The authors [28] introduced a novel approach to weakly supervised anomaly detection (WSVAD), reframing it as a supervised learning task with noisy labels, departing from the conventional MIL framework. Leveraging a Graph Convolutional Network (GCN), this method effectively cleans noisy labels, thereby enhancing the training and reliability of fully supervised action classifiers for anomaly detection. Additionally, the authors [57] introduced a binarization-embedded WSVAD (BE-WSVAD) method, which innovates by embedding binarization into GCN-based anomaly detection module.
@@
In [26], a novel approach called CLIPTSA leverages Vision Language (ViT) features for Weakly Supervised Video Anomaly Detection (WSVAD). Unlike conventional models such as C3D or I3D, CLIPTSA utilizes the Vision Transformer (ViT) encoded visual features from CLIP [34] to efficiently extract discriminative representations. It incorporates a Temporal Self-Attention (TSA) mechanism to model both long- and short-range temporal dependencies, thereby enhancing detection performance in VAD.
@@ V-C4 Unsupervised and Reconstruction-based Approaches
The reconstruction-based paradigm is typically categorized within certain learning frameworks as one-class classification (OCC) or unsupervised learning, primarily due to its emphasis on training using only the “normal” class of videos. Within the OCC paradigm, the model is trained only on data from one class (typically the “normal” class), and any deviation from this learned structure in the testing phase “anomaly” results in higher reconstruction error.
On the other hand, the unsupervised learning perspective focuses on understanding the structure or distribution within the data itself. Autoencoder-based anomaly detection is primarily unsupervised, as it learns to reconstruct data without explicitly being told what’s normal or anomalous. It just tries to minimize the difference between input and reconstructed output.
@@ V-F Anomaly Score
After calculating the PSNR between the predicted $\hat{Y}$ and its corresponding ground truth $Y$ for each frame $V_{i}$ in a test video and normalizing these values, we determine each frame’s regularity score $S_{r}(V_{i})$ using the equation below.
@@ VI-B Evaluation Metrics
Most previous work mainly compared their works on different datasets using the area under the ROC curve (AUC), Equal Error Rate (EER), and Average precision (AP) to measure their performance. The main aim of the metrics is to measure the model’s proficiency in differentiating between normal and abnormal videos. One significant metric used is the AUC “area under the Receiver Operating Characteristic (ROC) curve”, which plots the True Positive Rate (TPR) against the False Positive Rate (FPR) across various threshold settings. The TPR is defined as:
@@
$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}.$ (14)
A higher AUC value indicates a better model performance, where an AUC of 1 signifies a perfect model and an AUC of 0.5 suggests no discriminative power between positive and negative classes. The other method used is the AP, which computes the average value of precision for different recall values. It summarizes the precision-recall curve into a single value representing average performance across all threshold levels. Additionally, the EER represents the point at which the FPR equals the false negative rate (FNR). Lower EER values indicate a greater degree of accuracy in the system [56].
@@ VII-A Quantitative comparison
Method Venue Year Supervision type Feature extractor UCF-Crime AUC STech AUC XD-Violance AUC Avenue AUC Ref
Lu et al. ICCV 2013 Unsupervised 3D cube gradient and PCA 65.51 68.00 // // [15]
Hasan et al. CVPR 2016 Unsupervised Convolutional autoencoder 50.60 60.85 // 70.20 [16]
Sultani et al. CVPR 2018 Weakly Sup. C3D 75.41 // // // [11]
AMC ICCV 2019 Unsupervised Conv-AE // // // 86.9 [54]
MemAE ICCV 2019 Unsupervised 2D CNN Encoder-Decoder // 71.2 // 83.3 [68]
TCN-IBL ICIP 2019 Weakly Sup. C3D 78.66 82.50 // // [27]
GCN-TSN CVPR 2019 Weakly Sup. TSN - RGB 82.12 84.44 // // [28]
MPED-RNN CVPR 2019 Unsupervised RNN // 73.4 // // [69]
SRF IEEE SPL 2020 Weakly Sup. C3D 79.54 84.16 // // [29]
CVAE ICCV 2021 Unsupervised ML-MemAE-SC, CVAE // 76.2 // 91.1 [20]
SS-MTL CVPR 2021 self-supervised 3D CNN // // // 86.9 [9]
Noise Cleaner CVPRW 2021 Weakly Sup. C3D 78.26 84.16 // // [58]
DAM AVSS 2021 Weakly Sup. I3D- RGB 82.67 88.22 // // [30]
SA-CRF ICCV 2021 Weakly Sup. Relation-aware TSN ResNet-50 85.00 96.85 // // [31]
RTFM ICCV 2021 Weakly Sup. I3D- RGB 84.30 97.21 77.81 // [32]
MIST CVPR 2021 Weakly Sup. I3D- RGB 82.30 94.84 // // [47]
MSLNet-CTE AAAI 2022 Weakly Sup. VideoSwin-RGB 85.62 97.32 78.59 // [23]
GCL CVPR 2022 Unsupervised ResNext 71.04 // 78.93 // [67]
CUPL CVPR 2023 Weakly Sup. I3D+VGGish 86.22 // // // [70]
USTN-DSC CVPR 2023 Unsupervised Transformer Encoder-Decoder // 73.8 // 89.9 [50]
FPDM ICCV 2023 Unsupervised Diffusion Model 74.7 78.6 // // [71]
HSC CVPR 2023 Weakly Sup. AE // 83.4 // 93.7 [22]
SLMPT ICCV 2023 Unsupervised U-Net // 78.8 // 90.9 [66]
UMIL CVPR 2023 Weakly Sup. X-CLIP-B/32 86.75 // // // [62]
TeD-SPAD ICCV 2023 Weakly Sup. U-Net + I3D 75.06 // // // [72]
CLAV CVPR 2023 Weakly Sup. I3D-RGB 86.1 97.6 81.3 89.8 [64]
EVAL CVPR 2023 Unsupervised 3D CNN // 76.63 // 86.02 [73]
TEVAD CVPRW 2023 Weakly Sup. ResNet-50 I3D-RGB / SwinBERT 84.90 98.10 79.80 // [33]
CLIP-TSA ICIP 2023 Weakly Sup. CLIP 87.58 98.32 82.19 // [26]
@@
Another critical facet of the analysis pertains to performance metrics, particularly the Area Under the Curve (AUC) scores across different datasets, indicating notable enhancements in model performance over time. This trend not only signifies advancements in methodological efficacy but also underscores the increasing accuracy and reliability of anomaly detection systems. Leading models, notably recent ones like “CLIP-TSA” published in 2023, exhibit exceptionally high AUC scores, illustrating the rapid progress achieved in recent years, particularly with the integration of advanced techniques from Natural Language Processing (NLP) into models like CLIP, BLIP, and GPT.
@@ VIII Visualizing Bibliometric Networks for Thematic Analysis
TABLE III: Cluster central point and color for each cluster with representative paper and teheme

### v2

@@ V-B2 Deep Feature Extractors
While each feature representation has its strengths and weaknesses as empirically investigated by the SOTA VAD approaches [6, 26, 35, 55, 61]. The end-to-end deep feature learning paradigms are the recent trends and these approaches have advanced the VAD performance. We provide the following future directions in terms of the feature representations to improve detection, accuracy, and efficiency.
@@ V-C1 Supervised Approaches
In supervised learning, algorithms are developed using datasets where each frame is pre-labeled as either ’normal’ or ’anomalous.’ This allows the model to learn explicit distinctions between normal and abnormal events based on these annotations. However, the use of supervised methods in VAD is relatively rare. This is primarily because acquiring datasets with detailed, frame-level annotations is highly challenging and resource-intensive. Anomalous events are often rare and diverse, making it difficult to compile a comprehensive set of labeled anomalies. Additionally, the manual process of annotating each frame in extensive video data is laborious and time-consuming, further limiting the availability of such datasets. Consequently, while supervised learning can be powerful when precise labels are available, its application in VAD is constrained by the practical difficulties in obtaining extensive, accurately annotated training data.
@@ V-C2 Self-supervised Approaches
Self-supervised approaches involve training models using data that is not explicitly labeled for anomalies. Instead, the models learn to identify anomalous events by solving ”proxy tasks”, which generate supervisory signals from the data itself. These tasks are designed to be related to the main goal of detecting anomalies, helping the model to develop an understanding of normal patterns in the data without needing direct supervision from labeled examples of anomalous events. However, it demands careful design and selection of proxy tasks to ensure that the learned features are useful for identifying anomalous events. In the work by [9], researchers propose a method based on self-supervised and multi-task learning at the object level. Self-supervision involves training a 3D CNN on several proxy tasks that do not require labeled anomaly data. These tasks include determining the direction of object movement (arrow of time), identifying motion irregularities by comparing objects in consecutive versus intermittent frames, and reconstructing object appearances based on their preceding and succeeding frames. By learning from the video data itself what normal object behavior looks like, the model becomes adept at detecting anomalies when deviations from this learned behavior occur. This methodology enables effective anomaly detection even in the absence of explicit labels.
The authors continued their work in [63]. The new enhancements include integrating advanced object detection methods such as YOLOv5, optical flow, and background subtraction, which improve the detection of rapidly moving objects and those outside predefined classes. They also introduce transformer blocks into the architecture, exploring both 2D and 3D Convolutional Vision Transformers (CvT) to better capture complex spatiotemporal dependencies. These updates culminate in a more robust framework that significantly enhances the detection accuracy and adaptability in identifying anomalous events in video sequences.
@@ V-C3 Weakly Supervised Approaches
In weakly supervised approaches, unlike supervised approaches, obtaining precise frame-level annotations for anomalies in long video sequences can be challenging and time-consuming. Instead, annotators may label ”snippets” or short segments within the video where anomalies are observed, serving as weak labels for training the model.
The concept of weak supervision was first introduced by Sultani et al. [11] through their pioneering multiple instance learning (MIL) model. This model was the first to adopt the notion of weakly labeled training videos. In this approach, normal and anomalous videos are considered negative and positive bags, respectively, with video segments serving as instances in MIL. These bags are processed through feature extractors to capture spatiotemporal features before being directed to a fully connected network to produce the final output. The output anomaly score ranges from 0 to 1, with the goal of increasing the model’s output anomaly score for abnormal segments while decreasing it for normal segments. However, this method can introduce noisy labels since the weak annotations do not provide precise information about the exact location of anomalies within the segments. This ambiguity can lead to the model learning less accurate representations of normal and abnormal behaviors. Additionally, the reliance on segment-level labels can sometimes result in the inclusion of irrelevant frames within the labeled segments, further complicating the training process.
Other works have solved this issue. For example, the authors [28] introduced a novel approach to weakly supervised anomaly detection (WSVAD), reframing it as a supervised learning task with noisy labels, departing from the conventional MIL framework. Leveraging a Graph Convolutional Network (GCN), this method effectively cleans noisy labels, thereby enhancing the training and reliability of fully supervised action classifiers for anomaly detection. Additionally, the authors [64] introduced a binarization-embedded WSVAD (BE-WSVAD) method, which innovates by embedding binarization into GCN-based anomaly detection module.
@@
In [26], a novel approach called CLIPTSA leverages Vision Language (ViT) features for WSVAD. Unlike conventional models such as C3D or I3D, CLIPTSA utilizes the Vision Transformer (ViT) encoded visual features from CLIP [34] to efficiently extract discriminative representations. It incorporates a Temporal Self-Attention (TSA) mechanism to model both long- and short-range temporal dependencies, thereby enhancing detection performance in VAD.
Another recently published work by [61], proposes a novel weakly supervised framework called Text Prompt with Normality Guidance (TPWNG), leveraging the CLIP model for aligning textual descriptions with video frames to generate pseudo-labels. The method involves fine-tuning CLIP for domain adaptation using ranking and distributional inconsistency losses and introducing a learnable text prompt mechanism with normality visual prompts to improve text-video alignment. The framework also incorporates a pseudo-label generation module based on normality guidance to infer reliable frame-level pseudo-labels and a temporal context self-adaptive learning module to flexibly capture temporal dependencies in video events.
@@ V-C4 Unsupervised and Reconstruction-based Approaches
Although these models are feasible, scalable, and cost-effective approaches. However, the effectiveness of these models is highly dependent on the quality and comprehensiveness of the normal training data. If the normal data is not representative of all possible normal variations, the model’s performance in detecting anomalies can suffer.
Another issue is that the model could generate a high number of false positives, labeling normal variations or benign deviations as anomalies. This occurs because any deviation from the learned normal patterns, even if it is not truly anomalous, might be flagged.
@@
The evolution of the reconstruction approaches
The reconstruction-based paradigm is a key component of unsupervised learning frameworks and is often categorized within certain learning models as one-class classification (OCC) or unsupervised learning, primarily due to its emphasis on training using only the “normal” class of videos. Within the OCC paradigm, the model is trained only on data from one class (typically the “normal” class), and any deviation from this learned structure in the testing phase “anomaly” results in higher reconstruction error.
On the other hand, unsupervised learning focuses on understanding the structure or distribution within the data itself. It learns to reconstruct data without explicit labels indicating what is normal or anomalous.
As the work of, [68], proposed a novel approach called Generative Cooperative Learning, which combines a generator and a discriminator trained together using a negative learning paradigm. The generator, designed as an autoencoder, reconstructs both normal and abnormal representations. Through negative learning, the discriminator learns to estimate the probability of an instance being abnormal by identifying the reconstruction errors from the generator. This method leverages the assumption that anomalies are less frequent than normal events and that normal events exhibit temporal consistency, allowing for more effective anomaly detection.
Another unsupervised approach by [69] ) utilizes diffusion models, a type of generative model, to leverage the reconstruction capabilities for detecting anomalies. The approach begins by extracting features from video clips using a 3D convolutional neural network (3D-CNN). These features are then fed into a diffusion model, which reconstructs the features without relying on labeled data. The diffusion model progressively adds Gaussian noise to the input data and learns to reverse this process, effectively reconstructing the input.
@@ V-F Anomaly Score
After calculating the PSNR between the predicted $\hat{Y}$ and its corresponding ground truth $Y$ for each frame $V_{i}$ in a test video and normalizing these values, each frame’s regularity score is determined $S_{r}(V_{i})$ using the equation below.
@@ VI-B Evaluation Metrics
Most previous works mainly compared their results on different datasets using metrics such as the Area Under the ROC Curve (AUC) of the frame level, Equal Error Rate (EER), and Average Precision (AP) ) of the frame level to measure performance. The main aim of these metrics is to assess the model’s proficiency in differentiating between normal and abnormal videos.
Area Under the ROC Curve (AUC): The AUC is a significant metric used to evaluate model performance. It represents the area under the Receiver Operating Characteristic (ROC) curve, which plots the True Positive Rate (TPR) against the False Positive Rate (FPR) across various threshold settings. The TPR is defined as:
@@
$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$ (14)
A higher AUC value indicates better model performance, where an AUC of 1 signifies a perfect model and an AUC of 0.5 suggests no discriminative power between positive and negative classes. Two types of AUC that are commonly used for evaluating VAD models are Micro-AUC and Macro-AUC. Micro-AUC considers all samples across all classes together to compute a single AUC, making it sensitive to the most frequent classes. Macro-AUC, on the other hand, calculates the AUC for each class independently and averages these values, providing a balanced view of performance across all classes. While Micro-AUC reflects overall detection capability, Macro-AUC ensures the model performs well across different types of anomalies or normal events, offering insights into potential biases towards certain classes.
Average Precision (AP): The AP metric computes the average value of precision for different recall values. It summarizes the precision-recall curve into a single value representing average performance across all threshold levels.
Equal Error Rate (EER): The EER represents the point at which the FPR equals the False Negative Rate (FNR). Lower EER values indicate a greater degree of accuracy in the system [62].
@@ VII-A Quantitative comparison
Supervision type Method Venue Year Feature extractor UCF-Crime AUC STech AUC XD-Violance AUC Avenue AUC Ref
self-supervised SS-MTL CVPR 2021 3D CNN // // // 86.9 [9]
Unsupervised Lu et al. ICCV 2013 3D cube gradient and PCA 65.51 68.00 // // [15]
Hasan et al. CVPR 2016 Convolutional autoencoder 50.60 60.85 // 70.20 [16]
AMC ICCV 2019 Conv-AE // // // 86.9 [59]
MemAE ICCV 2019 2D CNN Encoder-Decoder // 71.2 // 83.3 [76]
MPED-RNN CVPR 2019 RNN // 73.4 // // [77]
CVAE ICCV 2021 ML-MemAE-SC, CVAE // 76.2 // 91.1 [20]
GCL CVPR 2022 ResNext 71.04 78.93 // // [68]
USTN-DSC CVPR 2023 Transformer Encoder-Decoder // 73.8 // 89.9 [55]
FPDM ICCV 2023 Diffusion Model 74.7 78.6 // // [78]
SLMPT ICCV 2023 U-Net // 78.8 // 90.9 [75]
EVAL CVPR 2023 3D CNN // 76.63 // 86.02 [44]
Weakly Sup. Sultani et al. CVPR 2018 C3D 75.41 // // // [11]
TCN-IBL ICIP 2019 C3D 78.66 82.50 // // [27]
GCN-TSN CVPR 2019 TSN - RGB 82.12 84.44 // // [28]
SRF IEEE SPL 2020 C3D 79.54 84.16 // // [29]
Noise Cleaner CVPRW 2021 C3D 78.26 84.16 // // [65]
DAM AVSS 2021 I3D- RGB 82.67 88.22 // // [30]
SA-CRF ICCV 2021 Relation-aware TSN ResNet-50 85.00 96.85 // // [31]
RTFM ICCV 2021 I3D- RGB 84.30 97.21 77.81 // [32]
MIST CVPR 2021 I3D- RGB 82.30 94.84 // // [52]
MSLNet-CTE AAAI 2022 VideoSwin-RGB 85.62 97.32 78.59 // [23]
CUPL CVPR 2023 I3D+VGGish 86.22 // // // [79]
HSC CVPR 2023 AE // 83.4 // 93.7 [22]
UMIL CVPR 2023 X-CLIP-B/32 86.75 // // // [71]
TeD-SPAD ICCV 2023 U-Net + I3D 75.06 // // // [80]
CLAV CVPR 2023 I3D-RGB 86.1 97.6 81.3 89.8 [73]
TEVAD CVPRW 2023 ResNet-50 I3D-RGB / SwinBERT 84.90 98.10 79.80 // [33]
CLIP-TSA ICIP 2023 CLIP 87.58 98.32 82.19 // [26]
TPWNG CVPR 2024 CLIP (ViT-B/16) 87.79 // 83.68 // [61]
@@
Another critical facet of the analysis pertains to performance metrics, particularly the Area Under the Curve (AUC) scores across different datasets, indicating notable enhancements in model performance over time. This trend not only signifies advancements in methodological efficacy but also underscores the increasing accuracy and reliability of anomaly detection systems. Leading models, notably recent ones like “CLIP-TSA” [26] published in 2023, exhibit exceptionally high AUC scores, illustrating the rapid progress achieved in recent years, particularly with the integration of advanced techniques from Natural Language Processing (NLP) into models like CLIP, BLIP, and GPT. A novel approach that utilized the vision language models is [61], published in 2024, which outperformed the CLIP-TSA over two datasets as shown in Table II. Based on the SOTA comparison, it is evident that Vision-language-based approaches are currently at the forefront of the VAD research.
@@ VIII Visualizing Bibliometric Networks for Thematic Analysis
TABLE III: Cluster central point and color for each cluster with representative paper and theme